When potassium chlorate is heated, it decomposes into potassium, chlorine and oxygen.

How many grams of potassium chlorate were used if 88

grams of oxygen, 56 grams of

potassium and 21 grams of chlorine were produced in the reaction?


1
Expert's answer
2021-03-26T05:42:30-0400

KClO3-> K+1/2Cl2 + 3/2O2

Molar Mass of KClO3 = 122.55g/mol

Molar Mass of K = 39.0983 g/mol

Molar Mass of O = 15.999g/mol

Molar Mass of Cl = 35.453 g/mol

= (56/39.0983) + (88/47.997) + (21/35.43)= 3.858

3.858×122.55 = 472.7979g